# The LHAASO Collaboration Measurement of attenuation length of the muon content in extensive air showers from 0.3 to 30 PeV with LHAASO

### v2

@@ I Introduction
The Pierre Auger Observatory has observed an excess of muon content in air showers with primary energy ranging from 6 to 16 EeV, which is higher than that predicted using LHC-tuned hadronic models [1]. This excess muon content within the same energy range was also confirmed by the Telescope Array (TA) experiment, which is the largest experiment conducted in the Northern Hemisphere [2]. Additionally, measurements conducted by the Sydney University Giant Air-shower Recorder (SUGAR), using a muon detector (MD), have shown an increasing excess of muons starting from 0.1 EeV and continuing with increasing primary energy [3]. In contrast, IceTop reported smaller muon densities than the post-LHC model predictions for energies ranging from 1 to 100 PeV [4]. The KASCADE-Grande observatory has also measured the attenuation length of the muon content of air showers at sea level for energies between 20 and 100 PeV and found that it is underestimated by all high-energy hadronic interaction models currently available [5]. Therefore, it can be concluded that none of the post-LHC high-energy hadronic interaction models can consistently describe the evolution of muon content in EASs in the aforementioned energy regime.
@@ II.1 LHAASO KM2A detector
LHAASO is a ground-based air shower observatory located 4410 m above sea level in Daocheng, China [6]. It contains a hybrid detector array comprising an EAS array covering an area of 1.3 km2 (KM2A), 78,000 m2 water Cherenkov detector array (WCDA), and 18 wide-field air Cherenkov/fluorescence telescopes (WFCTA). KM2A consists of 1188 MDs with a spacing of 30 m and 5216 electromagnetic detectors (EDs) with a spacing of 15 m; these detectors record muons and electromagnetic particles, respectively, as shown in Fig. 1.
@@
Each MD is composed of a water Cherenkov detector with an area of 36 m2. Fig. 2 shows a schematic of an MD. Each MD unit consists of a concrete tank with a diameter of 6.8 m and a depth of 1.2 m. The interior of the tank contains ultra-pure water enclosed in a highly reflective (95%) bag. Additionally, there is an overburden soil layer above the MD; it is approximately 2.5 m thick and absorbs electromagnetic and other charged particles in the shower. However, muons with energies greater than 1 GeV can penetrate overburden soil [7]. Thus, an 8-inch photo-multiplier tube (PMT) is positioned at the top surface of the water [8]. It is responsible for collecting the Cherenkov light emitted by high-energy muons. The total muon-sensitive area in LHAASO exceeds 40,000 m2. Moreover, owing to its wide dynamic range, the MDs can detect up to 10,000 particles, which enables accurate measurement of muon content across a broad energy range without saturation [6].
@@ II.2 Muon content measurement
When one muon particle passes through an MD, the Cherenkov light emitted from pure water is collected by the PMT after multiple reflections within the tank. The PMT signals pass through a shaping circuit and are then continuously digitized using flash analog-to-digital converters (FADCs) [6]. When the signal amplitude of one FADC channel exceeds the preset threshold, the signal is recorded as one ‘hit’ and this MD is recorded as a fired detector. An event trigger is generated if more than 20 EDs are fired within a 400 ns time window. In this work, we restrict the arrival time of particles to within [-30, 50] ns around the shower front, to suppress background or detector noise as doing in [9].
The total number of muons of a triggered event is the sum number of muons from all fired MDs. The number of muons of each fired MD are derived by dividing the number of photonelectrons (NPE) recorded by the MD by VEM/cos(). Here VEM is the vertical equivalent muon and is the zenith angle of the event. The VEM of each MD is calibrated once daily by almost vertical events [10]. An inclined muon travels a longer path in the tank water of the MD compared to a vertical muon, leading to a larger NPE. The NPE produced by a single muon is proportional to the path length; therefore, the equivalent muon for inclined showers should be VEM/cos() [7]. The distribution of muons from an individual MD in different directions is shown in the left-hand panel of Fig. 3. It can be observed that the single-muon and double-muon peaks are the same for each zenith angle bin. The right-hand panel of Fig. 3 shows the peak values of single-muon at each zenith angle range more clearly, and no zenith angle dependence was observed.
The high-energy electromagnetic particles near the core of the shower can punch through the shielding soil of an MD and become recorded as muon. To reduce the punch-through effect of high-energy electromagnetic particles near the shower core, hits within 40 m from the shower axis are not counted. Hits 200 m away from the shower axis are not counted either because of the limited size of KM2A. Therefore, in this study, the number of muons recorded by MDs within a distance of 40-200 m from the shower axis in a shower event was defined as the muon content of a shower event. According to [11], the muon content within this 40–200 m range exhibits a strong linear relationship with the true muon content.
@@ II.3 Monte Carlo simulation
Monte Carlo simulations were performed to compare the results with those obtained using interaction models. The COsmic Ray SImulations for KAscade (CORSIKA) code (version7.7410) [12] software package was used to simulate EASs of cosmic-rays with primary energies ranging from 100 TeV to 100 PeV. The zenith angles of the incident cosmic-rays were sampled within –. In CORSIKA simulation, the low-energy hadronic interaction model Fluka [13] and high-energy hadronic interaction models QGSJET-II-04 [14], EPOS-LHC [15], and SIBYLL 2.3d [16] were employed for model validation. The number of events simulated by the QGSJET-II-04 model exceeded 7.1107, whereas the number of events simulated by EPOS-LHC and SIBYLL 2.3d exceeded 3.5107. The primary particle components in the simulation included hydrogen (H), helium (He), nitrogen (N), aluminum (Al), and iron (Fe) nuclei.
The ED and MD detector responses were simulated using the G4KM2A [17] package developed in the framework of GEANT4 [18]. Random noise in a single ED and MD was also considered in the simulation.
The simulation data were normalized according to several popular assumed energy spectra, including Gaisser H3a [19], Horandel [20], and Global Spline Fit (GSF) [21]. A new spectrum model [22] developed according to the last published LHAASO spectrum results [23] was used; in this study, it is denoted as the LHAASO spectrum. The simulation results were validated using LHAASO KM2A data, which matched well with the data [9].
@@ II.4 Data and event selection
The atmospheric pressure has an impact on the measurement of muon content. By utilizing the relationship between the number of muons and the atmospheric pressure, we adjusted the number of muons to that under standard atmospheric pressure. The influence of atmospheric pressure on the measurement of muon content was 4.4, becoming 0.6 after correcting for atmospheric pressure [23].
The calibration uncertainty of one ED and MD for a single particle are less than 2 [10] and 0.5 [7], respectively. At 300 TeV, the average numbers of triggered EDs and MDs were approximately 200 and 50, respectively. Therefore, the measurement uncertainties of and are both less than 0.15. The zenith angle resolution is less than 0.2∘ at 400 TeV [9].
@@ III.1 Attenuation length measurement
The CIC method [24] was used to infer the attenuation length of muon content in air showers. This method is based on the assumption that the arrival distribution of cosmic-rays is isotropic; thus, the observed intensity of primary particles with the same energy is independent of the zenith angle or slant depth. This method depends on no supposed simulation models. The purpose of this study was to compare the muon content in air showers with the same primary energy at different zenith angles or different atmospheric depths.
@@ III.3 Comparison of experimental and simulated attenuation lengths
The LHAASO Collaboration has recently published the most accurate all-particle differential energy spectrum within the energy range of 0.3-30 PeV [23]. The all-particle energy spectrum of LHAASO fitted well with the following formula:
@@
The attenuation lengths obtained from the LHAASO data were shorter than those obtained from the simulation data, which were normalized to the LHAASO energy spectrum over the entire energy range. The results for the GSF energy spectrum were similar to those of the LHAASO energy spectrum. However, the attenuation lengths of the Gaisser H3a and Horandel energy spectra were shorter than those of the LHAASO and GSF spectra above 2 PeV, where the mean mass of the Gaisser H3a and Horandel models is heavier than that of the data [23].
The KASCADE-Grande Collaboration measured the attenuation length of the muon content for cosmic-ray air showers within 20-100 PeV at sea level. In particular, the muon content was measured by fitting the lateral distribution with a Lagutin-Raikin function with fixed shape. The attenuation length was 125685 gcm-2, which is longer than the results from simulation data [5]. The results from LHAASO data with similar energy were shorter than that from KASCADE data, as shown in Fig. 8, but close to the simulation results.

### v3

@@ I Introduction
The Pierre Auger Observatory has observed an excess of muon content in air showers with primary energy ranging from 6 to 16 EeV, which is higher than that predicted using LHC-tuned hadronic models [PAOtest]. This excess muon content within the same energy range was also confirmed by the Telescope Array (TA) experiment, which is the largest experiment conducted in the Northern Hemisphere [PhysRevD.98.022002]. Additionally, measurements conducted by the Sydney University Giant Air-shower Recorder (SUGAR), using a muon detector (MD), have shown an increasing excess of muons starting from 0.1 EeV and continuing with increasing primary energy [SUGAR]. In contrast, IceTop reported smaller muon densities than the post-LHC model predictions for energies ranging from 1 to 100 PeV [icetopden]. The KASCADE-Grande observatory has also measured the attenuation length of the muon content of air showers at sea level for energies between 20 and 100 PeV and found that it is underestimated by all high-energy hadronic interaction models currently available [KASCADEatt]. Therefore, it can be concluded that none of the post-LHC high-energy hadronic interaction models can consistently describe the evolution of muon content in EASs in the aforementioned energy regime.
@@ II.1 LHAASO KM2A detector
LHAASO is a ground-based air shower observatory located 4410 m above sea level in Daocheng, China [lhaaodesign]. It contains a hybrid detector array comprising an EAS array covering an area of 1.3 km2 (KM2A), 78,000 m2 water Cherenkov detector array (WCDA), and 18 wide-field air Cherenkov/fluorescence telescopes (WFCTA). KM2A consists of 1188 MDs with a spacing of 30 m and 5216 electromagnetic detectors (EDs) with a spacing of 15 m; these detectors record muons and electromagnetic particles, respectively, as shown in Fig. 1.
@@
Each MD is composed of a water Cherenkov detector with an area of 36 m2. Fig. 2 shows a schematic of an MD. Each MD unit consists of a concrete tank with a diameter of 6.8 m and a depth of 1.2 m. The interior of the tank contains ultra-pure water enclosed in a highly reflective (95%) bag. Additionally, there is an overburden soil layer above the MD; it is approximately 2.5 m thick and absorbs electromagnetic and other charged particles in the shower. However, muons with energies greater than 1 GeV can penetrate overburden soil [PRA2015_KM2Adesignzx]. Thus, an 8-inch photo-multiplier tube (PMT) is positioned at the top surface of the water [2022yupmt]. It is responsible for collecting the Cherenkov light emitted by high-energy muons. The total muon-sensitive area in LHAASO exceeds 40,000 m2. Moreover, owing to its wide dynamic range, the MDs can detect up to 10,000 particles, which enables accurate measurement of muon content across a broad energy range without saturation [lhaaodesign].
@@ II.2 Muon content measurement
When one muon particle passes through an MD, the Cherenkov light emitted from pure water is collected by the PMT after multiple reflections within the tank. The PMT signals pass through a shaping circuit and are then continuously digitized using flash analog-to-digital converters (FADCs) [lhaaodesign]. When the signal amplitude of one FADC channel exceeds the preset threshold, the signal is recorded as one ‘hit’ and this MD is recorded as a fired detector. An event trigger is generated if more than 20 EDs are fired within a 400 ns time window. In this work, we restrict the arrival time of particles to within [-30, 50] ns around the shower front, to suppress background or detector noise as doing in [cpclhaasocrab].
The total number of muons of a triggered event is the sum number of muons from all fired MDs. The number of muons of each fired MD are derived by dividing the number of photonelectrons (NPE) recorded by the MD by VEM/cos(). Here VEM is the vertical equivalent muon and is the zenith angle of the event. The VEM of each MD is calibrated once daily by almost vertical events [PRD2022EDcalLv]. An inclined muon travels a longer path in the tank water of the MD compared to a vertical muon, leading to a larger NPE. The NPE produced by a single muon is proportional to the path length; therefore, the equivalent muon for inclined showers should be VEM/cos() [PRA2015_KM2Adesignzx]. The distribution of muons from an individual MD in different directions is shown in the left-hand panel of Fig. 3. It can be observed that the single-muon and double-muon peaks are the same for each zenith angle bin. The right-hand panel of Fig. 3 shows the peak values of single-muon at each zenith angle range more clearly, and no zenith angle dependence was observed.
The high-energy electromagnetic particles near the core of the shower can punch through the shielding soil of an MD and become recorded as muon. To reduce the punch-through effect of high-energy electromagnetic particles near the shower core, hits within 40 m from the shower axis are not counted. Hits 200 m away from the shower axis are not counted either because of the limited size of KM2A. Therefore, in this study, the number of muons recorded by MDs within a distance of 40-200 m from the shower axis in a shower event was defined as the muon content of a shower event. According to [PRD2022zhyenergyreconstruction], the muon content within this 40–200 m range exhibits a strong linear relationship with the true muon content.
@@ II.3 Monte Carlo simulation
Monte Carlo simulations were performed to compare the results with those obtained using interaction models. The COsmic Ray SImulations for KAscade (CORSIKA) code (version7.7410) [corsikasim] software package was used to simulate EASs of cosmic-rays with primary energies ranging from 100 TeV to 100 PeV. The zenith angles of the incident cosmic-rays were sampled within –. In CORSIKA simulation, the low-energy hadronic interaction model Fluka [BATTISTONI201510] and high-energy hadronic interaction models QGSJET-II-04 [PhysRevD.83.014018], EPOS-LHC [PhysRevC.92.034906], and SIBYLL 2.3d [PhysRevD.102.063002] were employed for model validation. The number of events simulated by the QGSJET-II-04 model exceeded 7.1107, whereas the number of events simulated by EPOS-LHC and SIBYLL 2.3d exceeded 3.5107. The primary particle components in the simulation included hydrogen (H), helium (He), nitrogen (N), aluminum (Al), and iron (Fe) nuclei.
The ED and MD detector responses were simulated using the G4KM2A [km2asimulationnew] package developed in the framework of GEANT4 [g4toolkit]. Random noise in a single ED and MD was also considered in the simulation.
The simulation data were normalized according to several popular assumed energy spectra, including Gaisser H3a [Gaisser2013], Horandel [HORANDEL2003193], and Global Spline Fit (GSF) [ICRCGSF2017]. A new spectrum model [lv2024precise] developed according to the last published LHAASO spectrum results [PhysRevLett.132.131002] was used; in this study, it is denoted as the LHAASO spectrum. The simulation results were validated using LHAASO KM2A data, which matched well with the data [cpclhaasocrab].
@@ II.4 Data and event selection
The atmospheric pressure has an impact on the measurement of muon content. By utilizing the relationship between the number of muons and the atmospheric pressure, we adjusted the number of muons to that under standard atmospheric pressure. The influence of atmospheric pressure on the measurement of muon content was 4.4, becoming 0.6 after correcting for atmospheric pressure [PhysRevLett.132.131002].
The calibration uncertainty of one ED and MD for a single particle are less than 2 [PRD2022EDcalLv] and 0.5 [PRA2015_KM2Adesignzx], respectively. At 300 TeV, the average numbers of triggered EDs and MDs were approximately 200 and 50, respectively. Therefore, the measurement uncertainties of and are both less than 0.15. The zenith angle resolution is less than 0.2∘ at 400 TeV [cpclhaasocrab].
@@ III.1 Attenuation length measurement
The CIC method [CICmothed] was used to infer the attenuation length of muon content in air showers. This method is based on the assumption that the arrival distribution of cosmic-rays is isotropic; thus, the observed intensity of primary particles with the same energy is independent of the zenith angle or slant depth. This method depends on no supposed simulation models. The purpose of this study was to compare the muon content in air showers with the same primary energy at different zenith angles or different atmospheric depths.
@@ III.3 Comparison of experimental and simulated attenuation lengths
The LHAASO Collaboration has recently published the most accurate all-particle differential energy spectrum within the energy range of 0.3-30 PeV [PhysRevLett.132.131002]. The all-particle energy spectrum of LHAASO fitted well with the following formula:
@@
The attenuation lengths obtained from the LHAASO data were shorter than those obtained from the simulation data, which were normalized to the LHAASO energy spectrum over the entire energy range. The results for the GSF energy spectrum were similar to those of the LHAASO energy spectrum. However, the attenuation lengths of the Gaisser H3a and Horandel energy spectra were shorter than those of the LHAASO and GSF spectra above 2 PeV, where the mean mass of the Gaisser H3a and Horandel models is heavier than that of the data [PhysRevLett.132.131002].
The KASCADE-Grande Collaboration measured the attenuation length of the muon content for cosmic-ray air showers within 20-100 PeV at sea level. In particular, the muon content was measured by fitting the lateral distribution with a Lagutin-Raikin function with fixed shape. The attenuation length was 125685 gcm-2, which is longer than the results from simulation data [KASCADEatt]. The results from LHAASO data with similar energy were shorter than that from KASCADE data, as shown in Fig. 8, but close to the simulation results.